# STUPP-24-273

### v2

@@ 2.2 Lifetime
It should be remarked that in Fig. 2 (right) , the region with the escape time $T<10$ does not obey the power law. This is because this short-time behavior comes from the fact that no classical chaos occurred. These non-chaotic processes have been ignored in the fitting. In Fig. 2 (right) we also ignore the region with $T>200$ in the fitting because the sample number in this region is too little.
@@ 3.1 Reducing the BFSS matrix model
The BFSS matrix model is a $1+0$ dimensional $SU(N)$ gauge theory with adjoint matters. The bosonic part of the classical action is given by
@@
Each of $X^{r}$’s is an $N\times N$ hermitian matrix. The gauge field $A$ is also a hermitian matrix and non-dynamical. The variation with respect to $A$ leads to the Gauss law constraint:
@@ 3.2 Lifetime
$\tau=166.8\pm 0.1\,.$ (3.10)
This is the value for a certain set of the values of $E$ , $g$ , $l$ and $m$ The next task is to study the parameter dependence of the lifetime again.
@@ Acknowledgments
K. Y. is very grateful to Gordon Semenoff for the kind hospitality during his stay at UBC, where discussions were very beneficial in completing this research. The work of O. F. was supported by RIKEN Special Postdoctoral Researchers Program and JSPS Grant-in-Aid for Research Activity Start-up No. 24K22890. The work of T. S. was supported by JST SPRING, Grant Number JPMJSP2110. The work of K. Y. was supported by MEXT KAKENHI Grant-in-Aid for Transformative Research Areas A “Machine Learning Physics” No. 22H05115, and JSPS Grant-in-Aid for Scientific Research (B) No. 22H01217.

### v3

@@ 2.2 Lifetime
It should be remarked that in Fig. 2 (right) , the region with the escape time $T<10$ does not obey the power law. This is because this short-time behavior comes from the fact that no classical chaos occurred. These non-chaotic processes have been ignored in the fitting. In Fig. 2 (right) we also ignore the region with $T>150$ in the fitting because the sample number in this region is too little.
@@ 3.1 Reducing the BFSS matrix model
The BFSS matrix model is a $1+0$ dimensional $SU(N)$ gauge theory with adjoint matters [^2] The original gauge group is $U(N)$ , but the $U(1)$ is decoupled. Therefore, we take the group as $SU(N)$. The bosonic part of the classical action is given by
@@
Each of $X^{r}$’s is an $N\times N$ hermitian matrix and the action has the $SO(9)$ global $R$-symmetry. The constants $l_{s}$ and $g_{s}$ are string length and string coupling, respectively. The gauge field $A$ is also a hermitian matrix and non-dynamical. The variation with respect to $A$ leads to the Gauss law constraint:
@@ 3.2 Lifetime
$\tau=166.8\pm 0.1\,,$ (3.10)
for $E=0.1$ , $g=1$ , $l=5$ and $m=1$ The next task is to study the parameter dependence of the lifetime again.
@@ Acknowledgments
K. Y. is very grateful to Gordon Semenoff for the kind hospitality during his stay at UBC, where discussions were very beneficial in completing this research. The work of O. F. was supported by RIKEN Special Postdoctoral Researchers Program and JSPS Grant-in-Aid for Research Activity Start-up No. 24K22890. The work of T. S. was supported by JST SPRING, Grant Number JPMJSP2110. The work of K. Y. was supported by MEXT KAKENHI Grant-in-Aid for Transformative Research Areas A “Machine Learning Physics” No. 22H05115, and JSPS Grant-in-Aid for Scientific Research (B) No. 22H01217 and (C) No. 25K07313.
Appendix
Appendix A Parameters in the BFSS matrix model
Here, let us read off the relation among the parameters $m$ and $g$ in the reduced model (3.7) and the string theoretical parameters, $l_{s}$ (string length) and $g_{s}$ (string coupling).
The exact expression of the Lagrangian $L$ of the BFSS matrix model is given by
$L=\frac{1}{l_{s}g_{s}}{\rm tr}\left[\frac{1}{2}\sum_{r=1}^{9}(DX^{r})^{2}+\frac{1}{4(2\pi l_{s}^{2})^{2}}\sum_{r,s=1}^{9}[X^{r},X^{s}]^{2}\right]\,.$ (A.1)
Note here that the action (3.1) can be reproduced by setting $12\pi l_{s}^{2}=$ 1 . Then the Hamiltonian $H$ is given by
$H={\rm tr}\left[\frac{l_{s}g_{s}}{2}\sum_{r=1}^{9}(P_{r})^{2}-\frac{1}{16\pi^{2}l_{s}^{5}g_{s}}\sum_{r,s=1}^{9}[X^{r},X^{s}]^{2}\right]\,,$ (A.2)
where $P_{r}~{}(r=1,\ldots,9)$ are the canonical momenta.
By comparing the expression (A.2) with the reduced Hamiltonian (3.7) , we can read off the following relations:
$m=\frac{1}{l_{s}g_{s}}\,,\qquad g=\frac{1}{4\pi^{2}l_{s}^{5}g_{s}}\,.$ (A.3)
Then the lifetime is expressed as
$\tau\propto$ $\,E^{\alpha}\left(\frac{1}{l_{s}^{5}g_{s}}\right)^{-\alpha-\frac{1}{2}}\left(\frac{1}{l_{s}g_{s}}\right)^{\frac{1}{2}}l^{-4\alpha-1}=(g_{s}El_{s})^{\alpha}\left(\frac{l}{l_{s}}\right)^{-4\alpha-1}l_{s}\,.$ (A.4)
Note here that the cut-off $l$ remains as it is. Since our analysis is carried out at the classical level, $l$ is taken to be independent of $l_{s}$ and $g_{s}$. At the quantum mechanical level, it would of course depend on these, but this issue is beyond our scope here.